# Dephasing and error dynamics affecting a singlet-triplet qubit during coherent spin shuttling

### v1

@@ II.3 Repeated spin shuttles
To investigate dephasing during shuttling, we vary $t_{\mathrm{wait}}$ for a series of fixed $N$ carried out within $t_{\mathrm{shuttle}}$ (Fig. 3c), with $B_{0}\approx 0$ mT to minimize the charge noise-induced field-dependent noise dynamics on $\Delta E_{\mathrm{Z}}$ discussed above. We observe a single decay profile in singlet return, independent of $N$, indicating the decay is primarily time-dependent. A decay fit of the form in Eq. 3 to the combined data reveals $T_{2}^{*}=5.07\pm 0.07\leavevmode\nobreak\ \mu$s, which is similar to $T_{2}^{*}(\mathrm{transition})$ measured in the absence of shuttling and $B_{0}$ (Fig. 2c), and indicates the spin samples similar noise environments when shuttled and hybridized. This observation suggests the error introduced by the act of shuttling is low. The decay profile fits to a near-Gaussian with $\alpha=1.82\pm 0.07$, which points to the dominant dephasing mechanism being nuclear spin fluctuations.
Next, we employ a standard spin echo technique Jock2018 to reduce the effect of slow magnetic noise and extend the qubit dephasing time. This allows more shuttle operations to be performed within the qubit coherence time, to further highlight the role of shuttle-induced error dynamics. For this experiment, $B_{0}\approx 0$ mT to minimize field-dependent effects on $\Delta E_{\mathrm{Z}}$. We prepare the |$\ket{S}$ state and allow the two electron spins to evolve while one is stationary in QD0 and the other is repeatedly shuttled between QD1 and QD2 $N/2$ times. A calibrated $J$ pulse on B1 applies a $\pi_{z}$ operation to swap the two spins (Fig. 3b), followed by $N/2$ shuttles such that both spins sample the same magnetic environment for an equal amount of time, negating any acquired phase difference due to nuclear spin fluctuations slower than $1/2t_{\mathrm{shuttle}}$. Fig. 3d shows the singlet return probability after implementing the spin echo sequence during shuttling for a range of $N$ completed within $t_{\mathrm{shuttle}}$. We observe a single decay profile independent of $N$ and fit the data to a generic decay function of the form
@@
similar to Eq. 3. The fit yields $T_{2}^{\mathrm{Hahn}}=84.48\pm 7.82\leavevmode\nobreak\ \mu$s which is reasonably consistent with previously measured values in Si/SiGe heterostructures Kawakami2014 ; Kerckhoff2021 , and indicates sources of noise contributing to dephasing have been reduced. The fit converges with an exponent of $\alpha=0.64\pm 0.06$, which deviates from $\alpha$ commonly observed in the range of $1\lesssim\alpha\lesssim 2$ Petta2005 ; Kawakami2014 ; Veldhorst2014 ; Kobayashi2021 , suggesting a different noise spectrum is introduced. In other systems, sub-exponential decoherence behavior has been theorized to originate from 2D electron surfaces Witzel2012 , and has recently been predicted in 2D electron spin baths in NV- centers in diamond Schatzle2024 . The sub-exponential decay behavior observed here points to nontrivial noise dynamics of this nature affecting the qubit.
Next, we investigate the role of spin relaxation during shuttling since spin-flip mechanisms are known error sources in Si QDs Yang2013 . We prepare an eigenstate |$\ket{\uparrow\downarrow}_{(3,1,0)}$ of the $\Delta E_{\mathrm{Z}}$ interaction and monitor return of this state while repeatedly shuttling between QD1 and QD2. We used a small applied field of $B_{0}=10$ mT to avoid mixing with |$\ket{T_{-}}$ during the (3,1,0) charge state preparation. Fig. 3e shows singlet return probability of the shuttled $\ket{\uparrow\downarrow}$ state for a range of $N$ completed within $t_{\mathrm{shuttle}}$. We observe a single decay profile independent of $N$ and fit the data to an exponential decay function of the form
@@
with a range of assumptions placed on $y_{0}$, the singlet return saturation, since it is not explicitly measured. The fit results in $T_{1}=6.22\pm 0.71$ ms, where the uncertainty represents fit variation based on $y_{0}$, and lies within the broad range of $T_{1}$ measured for electrons confined to Si QDs, typically falling between milliseconds and seconds Simmons22011 ; Yang2013 ; Petit2018 ; Eng2015 . A variety of mechanisms generally stemming from spin-orbit and spin-valley interactions contribute to spin relaxation in Si QDs Tahan2014 ; Borjans2019 . Due to the natural disorder and inversion asymmetry at the Si/SiGe interface, the small in-plane $B_{0}$, as well as shuttling itself, give rise to spin-orbit effects Jock2018 that can induce spin flip relaxation over time. At $B_{0}=10$ mT, we expect a small but non-negligible influence from the spin-valley hotspot measured near 190 mT in QD0 (see discussion on Supplementary Fig. 4), which could drive a faster relaxation rate ($T_{1}^{-1}$) Yang2013 . Very few references for $T_{1}$ in the low-field limit exist; however, the observation that $T_{1}>T_{2}^{\mathrm{Hahn}}$ in our device shows that incoherent errors due to spin-flip processes during shuttling are relatively small compared to coherent errors arising from magnetic noise.
@@
From the data presented in Figs. 3c, d, and e, we anticipate the shuttling error to be low since the time-dependent singlet return decays in each panel are nearly indistinguishable among different $N$. To assess the time-independent error due to the shuttling operation, we fix $t_{\mathrm{shuttle}}$ while varying $N$ such that the effect of time-dependent dephasing and errors is constant with $N$. Figs. 3f, g, and h show the measured singlet return probability at three fixed $t_{\mathrm{shuttle}}$ corresponding to the shuttling experiments shown in the insets of Figs. 3c, d, and e, respectively. The error bars represent standard deviations of the mean of singlet return spanning repeated measurements totaling at least one hour. In all cases, the decay in singlet return is undetectable within error bars, and suggests shuttle errors are negligible out to at least $N=10^{3}$ (Fig. 3g). We make an attempt to quantify the worst-case estimate of shuttling error by fitting a simple linear function from the top of the first error bar to the bottom of the last error bar and extracting the slope of the line (not shown), which bounds the error per shuttle below $1\times 10^{-4}$, an improvement from recently reported shuttling errors vanRiggelen-Doelman2023 ; Struck2024 .
@@
Due to the low shuttling error, many repeated $N$ are required to identify and elucidate the nature of shuttle-induced losses. To do so we extend the previously discussed spin echo shuttling experiment out to the maximum $N$ possible within our instrument bandwidth capabilities. We compose a shuttling experiment at $B_{0}\approx 0$ mT to reduce field-dependent noise, with the minimum possible $42t_{\mathrm{wait}}=$ 4 ns and increase the repeated shuttles out to $N=4\times 10^{5}$, spanning a total time of $t_{\mathrm{shuttle}}=5.6$ ms, where $t_{\mathrm{shuttle}}(N)=14N$ ns. The singlet return probability over $N$, and equivalent $t_{\mathrm{shuttle}}$, is shown in Fig. 4. We consider the observed decay in the context of estimated $T_{2}^{\mathrm{Hahn}}$ and $T_{1}$ dynamics in this device, which are presented as shaded colored regions and are scaled by a multiplicative factor accounting for SPAM variation, serving as a guide to the eye. For a small number of shuttles $N<10^{3}$, the $T_{2}^{\mathrm{Hahn}}$ decay profile agrees with the data and implies the decay is dominated by the nontrivial spectral noise acting on $\Delta E_{\mathrm{Z}}$. For a large number of shuttles $N>10^{3}$, the decay deviates from this behavior and appears to be dominated by a different mechanism. The observation that the singlet return saturates below 50% indicates that the shuttling error is due to an incoherent spin relaxation process. The process causes decay on a timescale faster than the measured $T_{1}$ (for $N<10^{3}$), indicating a shuttle-induced, or $N$-dependent mechanism dominating spin relaxation when shuttling a large number of times ($N>10^{3}$).
@@
To characterize the nature of the shuttling error as a function of the number of shuttles, we fit to the data a phenomenological function of the form
@@
which defines the characteristic decay in units of $N$ ($N^{*}$). The fit results in a sub-exponential $\alpha=0.41\pm 0.03$, which suggests the presence of noise mechanisms that deviate from both a typical $1/f$-like noise spectrum observed for simple dephasing Petta2005 ; Connors2022 and a simple exponential decay expected of relaxation Simmons2011 ; Petit2018 . The observation that the function fits poorly for smaller $N<10^{3}$ suggests the influence of shuttle errors do not become apparent until after the completion of $N\sim 10^{3}$ shuttles. From the fit we also find $N^{*}=(1.92\pm 0.32)\times 10^{4}$ shuttles can be completed within a shuttle error-limited characteristic decay period, an order of magnitude more than recently demonstrated with shuttled spins Zwerver2023 ; vanRiggelen-Doelman2023 . Although a single decay profile is insufficient to describe the losses observed during shuttling, it points to the complexity of competing noise and error mechanisms at play during shuttling as $N$ increases.
@@ III Discussion
Understanding sources of error during shuttling is essential to the development of a successful spin shuttling platform used to entangle remote qubits. In this work, we have built on previous efforts exploring bucket brigade mode shuttling in Si Yoneda2021 ; Noiri2022 and in Ge vanRiggelen-Doelman2023 . Using a enhanced barrier pulse-assisted approach, we have demonstrated coherent spin shuttling in the limit of a large number of shuttles, with the goal of categorizing the different dephasing and relaxation mechanisms affecting the qubit while shuttling. Independent of shuttling, we identify magnetic noise from Overhauser field fluctuations as the dominant source of dephasing in the absence of external fields ($B_{0}\approx 0$ mT) and a combination of nuclear bath noise and charge noise originating from TLSs at the Si/SiGe interface at finite field ($B_{0}=50$ mT). When the spin is repeatedly shuttled, we observe little change in this decay at $B_{0}\approx 0$ mT, indicating that shuttling operations negligibly contribute to losses. Using a standard spin echo technique, we show coherence time during shuttling may be extended owing to the reduction of magnetic noise on $\Delta E_{\mathrm{Z}}$ and is limited by a different noise spectrum indicated by a subexponential $T_{2}^{\mathrm{Hahn}}$ decay profile. Residual unfiltered magnetic noise on $T_{2}^{\mathrm{Hahn}}$ timescales is determined to be the limiting time-dependent contribution since $T_{2}^{\mathrm{Hahn}}<T_{1}$. Additionally, we find that $T_{2}^{*}$, $T_{2}^{\mathrm{Hahn}}$, and $T_{1}$ are $N$-independent in the regime examined in this work and do not contribute appreciably to shuttling error.
By executing shuttle operations out to $N=4\times 10^{5}$, we are able to identify a shuttling-induced incoherent error which contributes significantly only after many shuttle operations. In the small $N$ ($<10^{3}$) regime, shuttling errors are negligible and decay is primarily limited by residual magnetic dephasing on $T_{2}^{\mathrm{Hahn}}$ timescales. In the large $N$ ($>10^{3}$) regime, a shuttle-induced, or $N$-dependent mechanism driving spin relaxation causes decay on a timescale faster than the measured $T_{1}$ (for $N<10^{3}$), which dominates and limits decay during the shuttling sequence. In the absence of temporal dephasing and relaxation (Figs. 3f., g., h.), we made a rudimentary yet encouraging shuttle error estimate of $1\times 10^{-4}$. A more thorough assessment is required to elucidate the details of this $N$-dependent process. Future work is required to improve characterization of the noise spectrum affecting the qubit during shuttling in order to extend shuttling capability, which can be done, for example, by spectrally targeting coherent noise on the gates Connors2022 .
@@
Lastly, we make estimates of the error per shuttle based on decay assumptions for 14 ns shuttling operations, summarized in Table 1. Of the dynamics described by the three error regimes evaluated in this work, error estimates based on the $T_{2}^{*}$ and $T_{1}$ decays represent the more practical expectation for building up a long-range spin shuttling platform. If fluctuating Overhauser fields are the only noise mechanism and decay is limited by $T_{2}^{*}$, we extrapolate the shuttling error to $2.2\times 10^{-5}$. In the case that spin-flip relaxation on the $T_{1}$ timescale limits decay, the error is estimated at $2.3\times 10^{-6}$. These results highlight the potential for low-error shuttling as a promising pathway to enabling distant spin entanglement in quantum computing platforms.

### v2

@@ II.3 Repeated spin shuttles
To investigate dephasing during shuttling, we vary $t_{\mathrm{wait}}$ for a series of fixed $N$ carried out within $t_{\mathrm{shuttle}}$ (Fig. 3c), with $B_{0}\approx 0$ mT to minimize the charge noise-induced field-dependent dynamics on $\Delta E_{\mathrm{Z}}$ discussed above. We observe a single decay profile in singlet return, independent of $N$, indicating the decay is primarily time-dependent. A decay fit of the form in Eq. 3 to the combined data reveals $T_{2}^{*}=5.07\pm 0.07\leavevmode\nobreak\ \mu$s, which is consistent with $T_{2}^{*}(\mathrm{transition})$ measured in the absence of shuttling at $B_{0}$ $\leavevmode\nobreak\ =0$ mT (Fig. 2c). This indicates the spin samples a similar noise environment when shuttled and hybridized. This observation suggests the error introduced by the act of shuttling is low. The decay profile is nearly Gaussian with $\alpha=1.82\pm 0.07$, which points to the dominant dephasing mechanism being nuclear spin fluctuations.
Next, we employ a standard spin echo technique Petta2005 to reduce the effect of slow magnetic noise and extend the qubit dephasing time. This allows more shuttle operations to be performed within the qubit coherence time, to further highlight the role of shuttle-induced errors. For this experiment, $B_{0}\approx 0$ mT to minimize field-dependent effects on $\Delta E_{\mathrm{Z}}$. We prepare the |$\ket{S}$ state and allow the two electron spins to evolve while one is stationary in QD0 and the other is repeatedly shuttled between QD1 and QD2 $N/2$ times. A calibrated $J$ pulse on B1 applies a $\pi_{z}$ operation to swap the two spins (Fig. 3b), followed by $N/2$ shuttles such that both spins sample the same magnetic environment for an equal amount of time, negating any acquired phase difference due to nuclear spin fluctuations slower than $1/2t_{\mathrm{shuttle}}$. Fig. 3d shows the singlet return probability after implementing the spin echo sequence during shuttling for a range of $N$ completed within $t_{\mathrm{shuttle}}$. We observe a single decay profile independent of $N$ and fit the data to a generic decay function of the form
@@
similar to Eq. 3. The fit yields $T_{2}^{\mathrm{Hahn}}=84.48\pm 7.82\leavevmode\nobreak\ \mu$s which is reasonably consistent with previously measured values in Si/SiGe heterostructures Kawakami2014 ; Kerckhoff2021 , and indicates sources of noise contributing to dephasing have been reduced. The fit converges with an exponent of $\alpha=0.64\pm 0.06$, which deviates from $\alpha$ commonly observed in the range of $1\lesssim\alpha\lesssim 2$ Petta2005 ; Kawakami2014 ; Veldhorst2014 , suggesting a different noise spectrum is introduced. Sub-exponential $T_{2}$ decays have been previously observed Stanwix2010 and simulated Schatzle2024 in other semiconductor-based spin systems, such as NV- centers in diamond, with origins tracing back to the particular spin dynamics of the bath. In our system, the exact mechanisms driving decay are as of yet unknown but may be related to bath dynamics at low magnetic field.
Next, we investigate the role of spin relaxation during shuttling since spin-flip mechanisms are known error sources in Si QDs Yang2013 . We prepare an eigenstate |$\ket{\uparrow\downarrow}_{(3,1,0)}$ of the $\Delta E_{\mathrm{Z}}$ interaction and monitor return of this state while repeatedly shuttling between QD1 and QD2. We use a small applied field of $B_{0}=10$ mT to avoid mixing with |$\ket{T_{-}}$ during the (3,1,0) charge state preparation. Fig. 3e shows singlet return probability of the shuttled $\ket{\uparrow\downarrow}$ state for a range of $N$ completed within $t_{\mathrm{shuttle}}$. We observe a single decay profile independent of $N$ and fit the data to an exponential decay function of the form
@@
with a range of assumptions placed on $y_{0}$, the singlet return saturation, since it is not explicitly measured. The fit results in $T_{1}=6.22\pm 0.71$ ms, where the uncertainty represents fit variation based on $y_{0}$, and lies within the broad range of $T_{1}$ measured for electrons confined to Si QDs, typically falling between milliseconds and seconds Petit2018 ; Eng2015 ; Borjans2019 . A variety of mechanisms generally stemming from spin-orbit and spin-valley interactions contribute to spin relaxation in Si QDs Tahan2014 . Due to the natural disorder and inversion asymmetry at the Si/SiGe interface, the small in-plane $B_{0}$ gives rise to spin-orbit effects Jock2018 that can induce spin-flip relaxation over time. At $B_{0}=0\sim 10$ mT, we expect a small but nonzero influence from the spin-valley hotspot measured near 190 mT in QD0, corresponding to a valley-splitting energy of $\sim$21 $\mu$eV in QD0 (see discussion on Supplementary Fig. 4), which could drive a faster relaxation rate ($T_{1}^{-1}$) Yang2013 . Spin-valley hotspots for the other QDs are higher in energy and would likely play a smaller role. Few references for $T_{1}$ in the low-field limit exist; however, the observation that $T_{1}>T_{2}^{\mathrm{Hahn}}$ in our device shows that incoherent errors due to spin-flip processes are relatively small compared to coherent errors arising from magnetic noise during shuttling for $N\lesssim 10^{3}$.
@@
From the data presented in Figs. 3c, d, and e, we anticipate the shuttling error to be low since the time-dependent singlet return decays in each panel are nearly indistinguishable among different $N$. To assess the time-independent error due to the shuttling operation, we fix $t_{\mathrm{shuttle}}$ while varying $N$ such that the effect of time-dependent dephasing and errors is constant with $N$. Figs. 3f, g, and h show the measured singlet return probability at three fixed $t_{\mathrm{shuttle}}$ corresponding to the shuttling experiments shown in the insets of Figs. 3c, d, and e, respectively. The error bars represent standard deviations of the mean of singlet return spanning repeated measurements totaling at least one hour. In all cases, the decay in singlet return is undetectable within error bars, and suggests shuttle errors are negligible out to at least $N\approx 10^{3}$ (Fig. 3g). We make an attempt to quantify the worst-case estimate of shuttling error by fitting a simple linear function from the top of the first error bar to the bottom of the last error bar and extracting the slope of the line (not shown), which bounds the error per shuttle below $10^{-4}$, an improvement from recently reported shuttling errors.
@@
Due to the low shuttling error, many repeated $N$ are required to identify and elucidate the nature of shuttle-induced losses. To do so we extend the previously discussed spin echo shuttling technique out to the maximum $N$ possible within our instrument bandwidth capabilities. We compose a shuttling experiment at $B_{0}\approx 0$ mT, with the minimum possible $42t_{\mathrm{wait}}=$ 4 ns and increase the repeated shuttles out to $N=4\times 10^{5}$, spanning a total time of $t_{\mathrm{shuttle}}=5.6$ ms, where $t_{\mathrm{shuttle}}(N)=14N$ ns. The singlet return probability over $N$, and equivalent $t_{\mathrm{shuttle}}$, is shown in Fig. 4. We consider the observed decay in the context of estimated $T_{2}^{\mathrm{Hahn}}$ and $T_{1}$ dynamics in this device, which are presented as shaded colored regions and are scaled by a multiplicative factor accounting for SPAM variation, serving as a guide to the eye. For a small number of shuttles $N<10^{3}$, the $T_{2}^{\mathrm{Hahn}}$ decay profile agrees with the data and implies the decay is dominated by the nontrivial magnetic spectral noise acting on $\Delta E_{\mathrm{Z}}$. For a large number of shuttles $N>10^{3}$, the decay deviates from this behavior and appears to be dominated by a different mechanism. The observation that the singlet return saturates below 50% indicates that the shuttling error is due to an incoherent spin relaxation process. However, the process causes decay on a timescale faster than the measured $T_{1}$ (for $N<10^{3}$), indicating a shuttle-induced, or $N$-dependent mechanism dominating spin relaxation when shuttling a large number of times ($N>10^{3}$).
Figure 4: Shuttling error estimates at large $N$. Data is shown in teal markers where error bars represent standard deviations of the mean of multiple repeated measurements. A phenomenological fit of the form of Eq. 7 is overlaid (black curve). The range of $T_{2}^{\mathrm{Hahn}}$ ($T_{1}$) based on analysis in the main text regarding Fig. 3d. (Fig. 3e.) is shown in orange (red) shading, presented as a guide to the eye. The top horizontal axis ($t_{\mathrm{shuttle}})$ is related to the bottom ($N$ shuttles) by the expression $t_{\mathrm{shuttle}}=N(2t_{\mathrm{wait}}+t_{\mathrm{ramp}})=14N\leavevmode\nobreak\ \mathrm{ns}$. Here, $B_{0}\approx 0$ mT.
@@
To characterize the nature of the shuttling error as a function of the number of shuttles, we fit the data to a phenomenological function of the form
@@
which defines the characteristic decay in units of $N$ ($N^{*}$). We find $N^{*}=(1.92\pm 0.32)\times 10^{4}$ shuttles can be completed within a shuttle error-limited characteristic decay period, 1-2 orders of magnitude more than recently demonstrated with shuttled spins vanRiggelen-Doelman2024 ; DeSmet2024 . The observation that the function fits poorly for smaller $N<10^{3}$ suggests the influence of shuttle errors do not become apparent until after the completion of $N\approx 10^{3}$ shuttles. The fit results in $\alpha=0.41\pm 0.03$ and, although a single decay profile is insufficient to describe the losses observed during shuttling, it points to the complexity of competing noise and error mechanisms at play during shuttling as $N$ increases.
@@ III Discussion
Understanding sources of error during shuttling is essential to the development of a successful spin shuttling platform used to entangle remote qubits. In this work, we have built on previous efforts exploring bucket brigade mode shuttling in Si Yoneda2021 ; Noiri2022 and in Ge vanRiggelen-Doelman2024 QDs. Using an enhanced barrier pulse-assisted approach, we have demonstrated coherent spin shuttling in the limit of a large number of shuttles, with the goal of categorizing the different dephasing and relaxation mechanisms affecting the qubit while shuttling. Independent of shuttling, we identify magnetic noise from Overhauser field fluctuations as the dominant source of dephasing in the absence of external fields ($B_{0}\approx 0$ mT) and a combination of nuclear bath noise and charge noise at finite field ($B_{0}=50$ mT). When the spin is repeatedly shuttled, we observe little change in the decay at $B_{0}\approx 0$ mT, indicating that shuttling operations marginally contribute to losses. Using a standard spin echo technique, we show the coherence time during shuttling may be extended owing to the reduction of magnetic noise on $\Delta E_{\mathrm{Z}}$ and is limited by a different noise spectrum related to hyperfine interactions indicated by a subexponential $T_{2}^{\mathrm{Hahn}}$ decay profile. Residual uncorrelated magnetic noise on $T_{2}^{\mathrm{Hahn}}$ timescales is determined to be the limiting time-dependent contribution since $T_{2}^{\mathrm{Hahn}}<T_{1}$. Additionally, we find that $T_{2}^{*}$, $T_{2}^{\mathrm{Hahn}}$, and $T_{1}$ appear to be $N$-independent in the small $N<10^{3}$ regime and do not contribute appreciably to shuttling error. Fitting the dephasing and relaxation data in this regime, (Figs. 3f., g., h.), we make a rudimentary yet encouraging shuttle error estimate of $10^{-4}$.
By executing shuttle operations in the large $N$ regime with a maximum of $N=4\times 10^{5}$ shuttles, we identify a shuttling-induced incoherent error which contributes significantly only after many shuttle operations. In the small $N$ ($<10^{3}$) regime, shuttling errors are negligible and decay is primarily limited by magnetic dephasing on $T_{2}^{\mathrm{Hahn}}$ timescales. In the large $N$ ($>10^{3}$) regime, a shuttle-induced, or $N$-dependent mechanism driving spin relaxation causes decay, which occurs on a timescale faster than the measured $T_{1}$ (for $N<10^{3}$). The exact mechanisms driving shuttle errors are unclear given the evidence, but their origins can be narrowed down. Nonadiabatic charge transfer is expected to contribute imperceptibly small errors due to the large tunnel coupling between QD1 and QD2, as previously noted. We also expect errors arising from electron-nuclear spin flip-flops when repeatedly ionizing QD1 and QD2 to be low ($<10^{-4}$) Witzel2022 .
On the other hand, the physical motion of the electron through the transition may give rise to additional relaxation channels. Charge noise is known to induce spin relaxation through driven transverse spin-orbit fluctuations, which would be most pronounced at the QD charge transition Borjans2019 . Furthermore, the strong charge hybridization at the transition could lead to an effective reduction in valley or orbital splitting as well as changes in the spin-valley coupling, which could enhance spin relaxation though spin-valley hotspot-like mechanisms Yang2013 . Reduced valley splitting or rapid change of valley phase when shuttling between QDs may also increase susceptibility to excitations outside of the valley ground state during transfer, but recent theoretical work suggests encouraging evidence that optimized control schedules may reduce this likelihood during spin shuttling Lima2024 . In the above scenarios, the probability of these spin-flip errors occurring would increase as more shuttles are appended, or as the shuttled electron spends more time near the charge transition. While these values have not been measured explicitly in the low magnetic field regime, they can be targeted as the focus of future studies expanding upon spin shuttling errors.
@@
Looking ahead, error estimates based on the $T_{2}^{*}$ and $T_{1}$ decays are the more practical expectation for building up a long-range spin shuttling platform. Focusing on the timescales quantified in this work ($T_{2}^{*}$ and $T_{1}$), we make estimates of the error per shuttle $p=1-e^{-(14\mathrm{ns}/\tau)^{\alpha}}$ based on the assumption of the limiting decay timescale $\tau$ during a 14 ns shuttling operation, summarized in Table 1. If all other noise sources were remediated and the fluctuating field gradient is the only remaining noise mechanism so that decay is limited by $T_{2}^{*}$, we extrapolate the shuttling error to be $p=2.2\times 10^{-5}$. If noise on $\Delta E_{\mathrm{Z}}$ could also be reduced to the level such that spin-flip relaxation on the $T_{1}$ timescale limits decay, the shuttle error is estimated even lower at $p=2.3\times 10^{-6}$. These results highlight the potential for high-fidelity shuttling as a promising pathway to remote spin entanglement in spin-based quantum computing platforms.